# Software paper for submission to the Journal of Open Research Software

### v1

@@ II. Implementation of the polynomial diffusion model
Firstly, we replicated some results from [17] using PDSim. Our replications precisely match the two figures from [17], as discussed in the section “Replicating Schwartz and Smith’s Results” on the GitHub page.
Secondly, we plotted the simulated and estimated futures prices with 95% confidence intervals for both the Schwartz and Smith model and the polynomial diffusion model. These plots specifically validate the implementation of the filtering methods. Users can reproduce these plots using the code provided in the sections “Tests for Schwartz and Smith Model” and “Tests for Polynomial Diffusion Model” on the GitHub page.
Finally, users can perform a unit test via the “Unit Tests” navigation bar of PDSim. This unit test calculates the coverage rate, defined as the proportion of simulated trajectories where over 95% of points fall within the confidence interval for a specific set of parameters. If the coverage rate exceeds 95%, we consider the set of parameters to be reasonable. A detailed introduction to the unit test is available in the “Unit Tests” section on the GitHub page.
@@
  • DT ($>=0.31$)
  • ggplot2 ($>=3.4.4$)
  • lubridate ($>=1.9.3$)
  • MASS ($>=7.3.60$)
  • plotly ($>=4.10.4$)
  • scales ($>=1.3.0$)
  • shiny ($>=1.8.0$)
  • shinythemes ($>=1.2.0$)
  • tidyr ($>=1.3.1$)
@@
Some of the functions in PDSim were originally developed and used in [10] to estimate hidden state variables and futures prices using the Schwartz-Smith model and the polynomial diffusion model. This package not only allows for detailed estimation but also offers a dynamic web application that vividly visualises the parameter identification challenges discussed in [10]. It serves as a valuable tool for researchers and practitioners in finance who are focused on commodity futures pricing. The interactive interface provides users with a deep understanding of both models, especially by illustrating the impact of individual parameter changes. The user-friendly design ensures accessibility for all users, even those with no programming experience. Additionally, the downloadable plots can be seamlessly incorporated into professional reports or academic papers.

### v2

@@ IV. Application to WTI crude oil futures data
First, we replicated selected results from [23] using PDSim. The reproduced figures match exactly with the two figures in [23], as presented in the “Replicating Schwartz and Smith’s Results” section on the GitHub page. Moreover, the mean absolute errors estimated by PDSim are very close to those reported in the original paper, as discussed in the “Application to WTI crude oil futures data” section of this paper.
Second, we plotted the simulated and estimated futures prices with 95% confidence intervals for both the Schwartz-Smith model and the polynomial diffusion model. These plots specifically validate the implementation of the filtering methods. Users can reproduce these plots using the code provided in the sections “Tests for Schwartz-Smith Model” and “Tests for Polynomial Diffusion Model” on the GitHub page.
Third, users can perform a unit test via the “Unit Tests” navigation bar of PDSim. This unit test calculates the coverage rate, defined as the proportion of simulated trajectories where over 95% of points fall within the confidence interval for a specific set of parameters. If the coverage rate exceeds 95%, we consider the set of parameters to be reasonable. A detailed introduction to the unit test is available in the “Unit Tests” section on the GitHub page.
Finally, we used the “tinytest” package to test the three main functions of PDSim, KF, EKF, and UKF. The outputs from PDSim are compared with corresponding analytical results. Users can re-run these tests after installing PDSim by executing:
tinytest::test_package(”PDSim”)
@@
  • DT (>=0.31>=0.31)
  • ggplot2 (>=3.4.4>=3.4.4)
  • lubridate (>=1.9.3>=1.9.3)
  • MASS (>=7.3.60>=7.3.60)
  • plotly (>=4.10.4>=4.10.4)
  • scales (>=1.3.0>=1.3.0)
  • shiny (>=1.8.0>=1.8.0)
  • shinythemes (>=1.2.0>=1.2.0)
  • tidyr (>=1.3.1>=1.3.1)
@@
Some of the functions in PDSim were originally developed and used in [15] to estimate latent state variables and futures prices using the Schwartz-Smith model and the polynomial diffusion model. This package not only allows for detailed estimation but also offers a dynamic web application that vividly visualises the parameter identification challenges discussed in [15]. It serves as a valuable tool for researchers and practitioners in finance who are focused on commodity futures pricing. The interactive interface provides users with a convenient tool for exploring both models, and analysing their sensitivity to individual parameter changes. The user-friendly design ensures accessibility for all users, even those with no programming experience. Additionally, the downloadable plots can be seamlessly incorporated into professional reports or academic papers.
(4) Discussions and Limitations
PDSim consists of both a Shiny application and an R package. The Shiny app provides a straightforward interface for simulating commodity futures data and estimating latent factors, making it accessible to a wide range of users. The R package, in contrast, contains the core functions and offers greater flexibility, allowing users to fit real-world data through R scripts. This separation is primarily due to computational considerations. Parameter estimation in high-dimensional spaces is computationally expensive, making R scripts more suitable than a web-based application for such tasks.
In financial modelling, stochastic processes with jumps are commonly employed. At present, the incorporation of jumps is out of the the scope of the current version of PDSim. However, the inclusion of jump components have been analysed in the following studies, [4], [8], or [21], and in particular, the polynomial diffusion model with jumps has been studied in [12] and its practical implementation would significantly broaden the applicability of PDSim. PDSim focuses on adapting and implementing the standard polynomial diffusion model rather than extending it to jump-diffusion settings.